Marshall Company is issuing eight-year bonds with a coupon rate of 5.56 percent and semiannual coupon payments. If the current market rate for similar bonds is 9.27 percent.


What will be the bond price? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and bond price to 2 decimal places, e.g. 15.25.)

Bond price $


If company management wants to raise $1.25 million, how many bonds does the firm have to sell? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and number of bonds to 0 decimal places, e.g. 5,275.)

Number of bonds bonds
Answer #1

Value of Bond = 1 1 (1+r) Coupon * + MaturityValue (1 + r) 7

r = 0.0927 /2 = 0.04635

n = 8 years * 2 = 16

Coupon = 0.0556 /2 0.0278

= \small 27.8 *\frac{1-\frac{1}{(1+0.04635)^{16}}}{0.04635} + \frac{1000}{(1+0.04635)^{16}}

= 308.272703255 + 484.36007928

= 793.63

Bond needs to be issued = 1,250,000 / 793.63 = 1575.04 OR 1576 Bonds

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.